# MonSter: Marry Monodepth to Stereo Unleashes Power

### v1

@@ Abstract
Stereo matching recovers depth from image correspondences. Existing methods struggle to handle ill-posed regions with limited matching cues, such as occlusions and textureless areas. To address this, we propose MonSter, a novel method that leverages the complementary strengths of monocular depth estimation and stereo matching. MonSter integrates monocular depth and stereo matching into a dual-branch architecture to iteratively improve each other. Confidence-based guidance adaptively selects reliable stereo cues for monodepth scale-shift recovery. The refined monodepth is in turn guides stereo effectively at ill-posed regions. Such iterative mutual enhancement enables MonSter to evolve monodepth priors from coarse object-level structures to pixel-level geometry, fully unlocking the potential of stereo matching. As shown in Fig. 2, MonSter ranks $1^{st}$ across five most commonly used leaderboards — SceneFlow, KITTI 2012, KITTI 2015, Middlebury, and ETH3D. Achieving up to 49.5% improvements (Bad 1.0 on ETH3D) over the previous best method. Comprehensive analysis verifies the effectiveness of MonSter in ill-posed regions. In terms of zero-shot generalization, MonSter significantly and consistently outperforms state-of-the-art across the board. The code is publicly available at: https://github.com/Junda24/MonSter.
{strip}[Uncaptioned image]
Figure 1: Zero-shot generalization comparison: all models are trained on Scene Flow and tested directly on KITTI. Compared to the baseline IGEV [38], our method MonSter shows significant improvement in challenging regions such as reflective surfaces, textureless areas, fine structures, and distant objects.
1 Introduction
Figure 2: Leaderboard performance. Our method ranks $1^{st}$ across 5 leaderboards, advancing SOTA by a large margin.
Stereo matching estimates a disparity map from rectified stereo images, which can be subsequently converted into metric depth. It is the core of many applications such as self-driving, robotic navigation, and 3D reconstruction.
@@ 1 Introduction
Deep learning based methods [4, 14, 48, 20, 16, 35] have demonstrated promising performance on standard benchmarks. These methods can be roughly categorized into cost filtering-based methods and iterative optimization-based methods. Cost filtering-based methods [4, 14, 37, 15, 41, 22, 43, 39] construct 3D/4D cost volume using CNN features, followed by a series of 2D/3D convolutions for regularization and filtering to minimize mismatches. Iterative optimization-based methods [20, 16, 38, 49, 10] initially construct an all-pairs correlation volume, then index a local cost to extract motion features, which guide the recurrent units (ConvGRUs) [9] to iteratively refine the disparity map.
@@
Both types of methods essentially derive disparity from similarity matching, assuming visible correspondences in both images. This poses challenges in ill-posed regions with limited matching cues, e.g., occlusions, textureless areas, repetitive/thin structures, and distant objects with low pixel representation. Existing methods [4, 37, 44, 31, 42, 33, 49] address this issue by enhancing matching with stronger feature representations during feature extraction or cost aggregation: GMStereo [42] and STTR [18] employ transformer as feature extractors; IGEV [38] and ACVNet [37] incorporate geometric information into the cost volume using attention mechanisms to strengthen matching information; Selective-IGEV [35] and DLNR [49] further improve performance by introducing high-frequency information during the iterative refinement stage. However, these methods still struggle to fundamentally resolve the issue of mismatching, limiting their practical performance.
Figure 3: Distance between GT disparity and the disparity derived from monocular depth [46] on KITTI dataset. The red line indicates identical disparity maps. (a): Without any alignment. (b) Align depth with GT using global scale and global shift values (same for all pixels). (c) The monocular disparity produced by MonSter, with per-pixel shift refinement. Even globally aligned with GT, SOTA monocular depth models still exhibit severe noise. Our method MonSter effectively addresses this issue by refining monocular depth with per-pixel shift, which fully unlocks the power of monocular depth priors for stereo matching.
@@
Unlike stereo matching, monocular depth estimation directly recovers 3D from a single image, which does not encounter the challenge of mismatching. While monocular depth provides complementary structural information for stereo, pre-trained models often yield relative depth with scale and shift ambiguities. As shown in Fig.3, the prediction of monodepth models differ heavily from the ground-truth. Even after global scale and shift alignment, substantial errors still persist, complicating pixel-wise fusion of monocular depth and stereo disparity. Based on these insights, we propose MonSter, a novel approach that decomposes stereo matching into monocular depth estimation and per-pixel scale-shift recovery, which fully combines the strengths of monocular and stereo algorithms and overcomes the limitations from the lack of matching cues.
@@
MonSter constructs separate branches for monocular depth estimation and stereo matching, and adaptively fuses them through Stereo Guided Alignment (SGA) and Mono Guided Refinement (MGR) modules. SGA first rescales monodepth into a “monocular disparity” by aligning globally with stereo disparity. Then it uses condition-guided GRUs to adaptively select reliable stereo cues for updating the per-pixel monocular disparity shift. Symmetric to SGA, MGR uses the optimized monocular disparity as the condition to adaptively refine the stereo disparity in regions where matching fails. Through multiple iterations, the two branches effectively complement each other: 1) Though beneficial at coarse object-level, directly and unidirectionally fuse monodepth into stereo suffers from scale-shift ambiguities, which often introduces noise in complex regions such as slanted or curved surfaces. Refining monodepth with stereo resolves this issue effectively, ensuring the robustness of MonSter. 2) The refined monodepth in turn provides strong guidance to stereo in challenging regions. E.g., the depth perception ability of stereo matching degrades with distance due to smaller pixel proportions and the increased matching difficulty. Monodepth models pre-trained on large-scale datasets are less affected by such issues, which can effectively improve stereo disparity in the corresponding region.
Our main contributions can be summarized as follows:
  • We propose a novel stereo-matching method MonSter, fully leveraging the pixel-level monocular depth priors to significantly improve the depth perception performance of stereo matching in ill-posed regions and fine structures.
  • As shown in Fig. 2, MonSter ranks $1^{st}$ across 5 widely-used leaderboards: KITTI 2012 [11], KITTI 2015 [23], Scene Flow [22], Middlebury [27] and ETH3D [28]. Advancing SOTA by up to $49.5\%$.
  • Compared to SOTA methods, MonSter achieves the best zero-shot generalization consistently across diverse datasets. MonSter trained solely on synthetic data demonstrates strong performance across diverse real-world datasets (see Fig.1).
2 Related Work
@@
Matching based Methods. Mainstream stereo matching methods recover disparity from matching costs. These methods can generally be divided into two categories: cost filtering-based methods and iterative optimization-based methods. Cost filtering-based methods [14, 22, 43, 6, 5, 12, 29] typically construct a 3D/4D cost volume using feature maps and subsequently employing 2D/3D CNN to filter the volume and derive the final disparity map. Constructing a cost volume with strong representational capacity and accurately regressing disparity from a noisy cost volume are key challenges. PSMNet [4] proposes a stacked hourglass 3D CNN for better cost regularization. [14, 37, 30] propose the group-wise correlation volume, the attention concatenation volume, and the pyramid warping volume respectively to improve the expressiveness of the matching cost. Recently, a novel class of methods based on iterative optimization [20, 16, 38, 49, 35] has achieved state-of-the-art performance in both accuracy and efficiency. These methods use ConvGRUs to iteratively update the disparity by leveraging local cost values retrieved from the all-pairs correlation volume. Similarly, iterative optimization-based methods also primarily focus on improving the matching cost construction and the iterative optimization stages. CREStereo [16] proposes a cascaded recurrent network to update the disparity field in a coarse-to-fine manner. IGEV [38] proposes a Geometry Encoding Volume that encodes geometry and context information for a more robust matching cost. Both types of methods essentially derive depth from matching costs, which are inherently limited by ill-posed regions.
Stereo Matching with Structural Priors. Matching in ill-posed regions is challenging, previous methods [44, 31, 47, 7, 19, 17] tried to leverage structural priors to address this issue. EdgeStereo [31] enhances performance in edge regions by incorporating edge detection cues into the disparity estimation pipeline. SegStereo [44] utilizes semantic cues from a segmentation network as guidance for stereo matching, improving performance in textureless regions. However, semantic and edge cues only provide object-level priors, which are insufficient for pixel-level depth perception. Consequently, they are not effective in challenging scenes with large curved or slanted surfaces. Therefore, some methods incorporate monocular depth priors to provide per-pixel guidance for stereo matching through dense relative depth. CLStereo [47] introduces a monocular branch that serves as a contextual constraint, transferring geometric priors from the monocular to the stereo branch. Los [17] uses monocular depth as local structural prior to generating the slant plane, which can explicitly leverage structure information for updating disparities. However, monocular depth estimation suffers from severe scale and shift ambiguities as shown in Fig.3. Directly using it as a structural prior to constrain stereo matching can introduce heavy noise. Our MonSter adaptively selects reliable stereo disparity to correct the scale and shift of monodepth, which fully leverages the monocular depth priors while avoiding noise, thereby significantly enhancing stereo performance in ill-posed regions.
3 Method

### v2

# MonSter++: Unified Stereo Matching, Multi-view Stereo, and Real-time Stereo with Monodepth Priors

@@ Abstract
We introduce MonSter++, a geometric foundation model for multi-view depth estimation, unifying rectified stereo matching and unrectified multi-view stereo. Both tasks fundamentally recover metric depth from correspondence search and consequently face the same dilemma: struggling to handle ill-posed regions with limited matching cues. To address this, we propose MonSter++, a novel method that integrates monocular depth priors into multi-view depth estimation, effectively combining the complementary strengths of single-view and multi-view cues. MonSter++ fuses monocular depth and multi-view depth into a dual-branched architecture. Confidence-based guidance adaptively selects reliable multi-view cues to correct scale ambiguity in monocular depth. The refined monocular predictions, in turn, effectively guide multi-view estimation in ill-posed regions. This iterative mutual enhancement enables MonSter++ to evolve coarse object-level monocular priors into fine-grained, pixel-level geometry, fully unlocking the potential of multi-view depth estimation. MonSter++ achieves new state-of-the-art on both stereo matching and multi-view stereo. By effectively incorporating monocular priors through our cascaded search and multi-scale depth fusion strategy, our real-time variant RT-MonSter++ also outperforms previous real-time methods by a large margin. As shown in Fig. 1, MonSter++ achieves significant improvements over previous methods across eight benchmarks from three tasks—stereo matching, real-time stereo matching, and multi-view stereo, demonstrating the strong generality of our framework. Besides high accuracy, MonSter++ also demonstrates superior zero-shot generalization capability. We will release both the large and the real-time models to facilitate their use by the open-source community. The code will be released at: https://github.com/Junda24/MonSter-plusplus.
I Introduction
Understanding 3D geometry from 2D images is a fundamental problem in computer vision and serves as the cornerstone of numerous real-world applications, such as autonomous driving, robotic navigation, and augmented reality. Among the many facets of 3D perception, absolute depth estimation plays a pivotal role, as it provides metric scene understanding that is crucial for downstream tasks involving interaction, planning, and decision-making.
Figure 1: The proposed MonSter++ vs previous SOTA on stereo matching, multi-view stereo (mvs), and real-time stereo. MonSter++ demonstrates strong multi-task generalization, advanced previous SOTA by a large margin in all three tasks, and ranks $1^{st}$ across 8 leaderboards.
Obtaining accurate metric depth from multi-view data such as rectified stereo matching [1, 2, 3, 4, 5, 6, 7] and unrectified multi-view stereo [8, 9, 10, 11, 12, 13] has been widely adopted in practice. Both methods recover depth by establishing dense correspondences across multiple views along epipolar lines. In stereo matching, rectified image pairs constrain the epipolar lines to be horizontal, simplifying the correspondence search. In contrast, multi-view stereo handles images with arbitrary camera poses, where matching is performed along more general epipolar directions. In this paper, we adopt the term multi-view depth estimation as a unified formulation that encompasses both traditional stereo matching (two-view setting) and more general multi-view setups (three or more views).
@@ I Introduction
With the rapid advancement of deep learning and large-scale datasets, learning-based methods [1, 2, 4, 6, 5, 7, 8, 9, 14, 10, 11, 12, 13, 15, 16, 17, 18, 19] have dominated multi-view depth estimation and the corresponding leaderboards [20, 21, 22, 23, 24, 25]. Mainstream multi-view depth estimation methods typically follow a four-stage pipeline: feature extraction, cost volume construction, cost aggregation, and depth regression. The most critical stages are cost volume construction and cost aggregation. While cost volume construction offers initial similarity estimates for potential correspondences across views, it is often prone to noise and ambiguity. Therefore, effective cost aggregation and regularization are crucial for refining the volume and achieving accurate, robust depth estimation. Based on different cost aggregation strategies, these methods can be roughly categorized into cost filtering-based methods and iterative optimization-based methods. Cost filtering-based methods [1, 2, 26, 27, 28, 22, 29] construct 3D/4D cost volume using CNN features, followed by a series of 2D/3D convolutions for regularization and filtering to minimize mismatches. Iterative optimization-based methods [4, 30, 6, 31] initially construct an all-pairs correlation volume, then index a local cost to extract motion features, which guide the recurrent units (ConvGRUs) [32] to iteratively refine the depth map.
Figure 2: Zero-shot generalization comparison on stereo matching benchmarks. To validate the effectiveness of our method, we compare both our large model MonSter++ and the real-time variant RT-MonSter++ against state-of-the-art methods. Our method achieves particularly notable improvements on fine object boundaries in both accuracy models and real-time methods. Remarkably, RT-MonSter++ requires only 47 ms per inference yet even surpasses the accuracy-focused SOTA method IGEV++ [33], which takes 280 ms, particularly on edge regions. This highlights the superior efficiency of our method.
@@
Both types of methods essentially derive depth through similarity matching, assuming the presence of visible correspondences across images. This poses challenges in ill-posed regions with limited matching cues, e.g., occlusions, textureless areas, repetitive/thin structures, and distant objects with low pixel representation. Existing methods [1, 26, 34, 35, 29, 36, 31, 37, 12, 15] address this issue by enhancing matching with stronger feature representations during feature extraction or cost aggregation. However, these methods still struggle to fundamentally resolve the issue of mismatching, limiting their practical performance.
@@
Unlike multi-view depth estimation, monocular depth estimation directly recovers 3D from a single image, thereby avoiding the challenge of mismatching in ill-posed regions. While monocular depth provides complementary structural information for stereo, pre-trained models often yield relative depth with scale and shift ambiguities. As shown in Fig.3, the prediction of monodepth models differs significantly from the ground truth. Even after global scale and shift alignment, substantial errors still persist, complicating pixel-wise fusion of monocular depth and multi-view depth. In this work, we formulate multi-view depth estimation as a problem of leveraging multi-view matching information to achieve per-pixel scale-shift recovery based on monocular depth estimation. We argue that this approach effectively combines the advantages of both single-view and multi-view algorithms while overcoming the limitations caused by the lack of matching cues.
Figure 3: Distance between GT disparity and the disparity derived from monocular depth [38] on KITTI dataset. The red line indicates identical disparity maps. (a): Without any alignment. (b) Align depth with GT using global scale and global shift values (same for all pixels). (c) The monocular disparity produced by MonSter++, with per-pixel shift refinement. Even globally aligned with GT, SOTA monocular depth models still exhibit severe noise. Our method MonSter++ effectively addresses this issue by refining monocular depth with per-pixel shift, which fully unlocks the power of monocular depth priors for stereo matching.
@@
To this end, we propose MonSter++, which constructs separate branches for single-view and multi-view depth estimation, and adaptively fuses them through Stereo Guided Alignment (SGA) and Mono Guided Refinement (MGR) modules. SGA initially rescales the monocular depth to a coarse metric scale via global alignment with multi-view depth. Then it uses condition-guided GRUs to adaptively select reliable stereo cues for updating the per-pixel monocular depth shift. Symmetric to SGA, MGR uses the optimized monocular depth as the condition to adaptively refine the multi-view depth in regions where matching fails. Through multiple iterations, the two branches effectively complement each other: 1) Though beneficial at coarse object-level, directly and unidirectionally fusing monodepth into multi-view stereo suffers from scale-shift ambiguities, which often introduces noise in complex regions such as slanted or curved surfaces. Refining monodepth with stereo resolves this issue effectively, ensuring the robustness of MonSter++. 2) The refined monodepth in turn provides strong guidance to multi-view depth in challenging regions. E.g., the perception ability of multi-view depth degrades with distance due to smaller pixel proportions and the increased matching difficulty. Monodepth models pre-trained on large-scale datasets are less affected by such issues, which can effectively improve multi-view depth in the corresponding region.
By fully leveraging the strengths of monocular depth priors, MonSter++ achieves state-of-the-art performance in both stereo matching and multi-view stereo, ranking first on four most commonly used stereo matching benchmarks  [20, 21, 39, 23] and leading two key multi-view depth benchmarks [24, 25]. In addition to improving accuracy, MonSter++ substantially enhances the generalization ability of stereo matching, achieving competitive results on real-world scenarios with minimal synthetic training data and surpassing existing state-of-the-art methods. To fully leverage the advantages of our depth fusion architecture, we also introduce a real-time version of MonSter++, named RT-MonSter++, offering an attractive solution for time-constrained applications. By leveraging monocular priors as in MonSter++ and adopting an efficient cascaded search and multi-scale depth fusion strategy, RT-MonSter++ sets a new state of the art among real-time methods on the KITTI benchmarks, while also delivering strong generalization. Notably, RT-MonSter++ even surpasses the accuracy-oriented SOTA method IGEV++, particularly on thin structures, as shown in Fig. 2. Its balance of accuracy and speed makes it highly practical for deployment on mobile platforms like drones, autonomous vehicles, and robotic systems.
Figure 4: Zero-shot generalization comparison: all models are trained on Scene Flow and tested directly on KITTI. Compared to the baseline IGEV [6], our method MonSter++ shows significant improvement in challenging regions such as reflective surfaces, textureless areas, fine structures, and distant objects.
Experimental results show that our method significantly outperforms previous SOTA methods across all three tasks—stereo matching, multi-view stereo, and real-time stereo matching—highlighting the efficiency and versatility of our framework. As shown in Fig. 1, MonSter++ ranks $1^{st}$ across 4 widely-used stereo matching leaderboards: KITTI 2012 [39], KITTI 2015 [23], Middlebury [21] and ETH3D [20]. It advances SOTA by up to $21.8\%$ compared to the previous best method on KITTI 2012. As for multi-view stereo task, MonSter++ also ranks $1^{st}$ across 2 widely-used multi-view depth benchmarks: DDAD [24] and KITTI Eigen split [25]. By leveraging a coarse-to-fine framework, we extend MonSter++ into a more efficient and lightweight variant, RT-MonSter++, which achieves real-time inference while surpassing all existing real-time methods on the KITTI benchmark in accuracy. RT-MonSter++ outperforms the previous best real-time method RT-IGEV by 20.0% on the KITTI 2012 benchmark. In addition, by effectively leveraging monocular priors, both MonSter++ and RT-MonSter++ exhibit superior zero-shot generalization. Under the same training setup, our methods achieve the highest zero-shot performance on KITTI 2012, KITTI 2015, ETH3D, and Middlebury datasets. Notably, RT-MonSter++ requires only 47 ms for inference to reach performance comparable to the accuracy-oriented model IGEV (180 ms), and even surpasses it on three out of the four datasets. Our contributions can be summarized below:
  • We propose a unified framework MonSter++ for multi-view depth estimation that handles both rectified and unrectified input images. MonSter++ fully leverages the pixel-level monocular depth priors to significantly improve the accuracy in ill-posed regions and fine structures. As shown in Fig. 1, MonSter++ ranks $1^{st}$ across 4 widely-used stereo matching leaderboards and 2 mainstream multi-view stereo benchmarks.
  • We propose RT-MonSter++, which achieves real-time inference while delivers the best accuracy among all published real-time methods on the KITTI 2012 [39] and 2015 [23] benchmarks.
  • We perform large scale training on 2M image pairs collected from public datasets, enabling superior zero-shot generalization over previous methods consistently across diverse datasets for both MonSter++ and RT-MonSter++. All models and training code will be released to the community.
II Related Work
Depth estimation is a key technology in 3D perception, crucial for downstream tasks such as 3D reconstruction [11, 13, 12, 40, 15], localization [41, 42, 43, 44, 45, 46, 47, 48, 49, 50], and decision-making. Existing visual depth estimation methods can be broadly categorized into monocular depth estimation, stereo matching, and multi-view stereo. We first review related work of each task independently, and then discuss how recent studies unify them by incorporating monocular depth priors and/or semantic cues into multi-view depth estimation.
II-A Monocular depth estimation
Monocular depth estimation is a fundamental problem in computer vision. In recent years, CNN-based methods [51, 52, 53, 54, 55, 56] have gained considerable attention, typically formulating the task as either a per-pixel classification [52, 57, 54] or regression problem [58, 59, 60, 61]. To enhance performance, some approaches focus on improving feature representations [53, 62], while others propose more effective loss functions [52, 51, 54] or utilize mixed-data training strategies [63, 64, 65]. Recent monocular depth estimation methods [66, 38, 67, 68], built upon powerful image backbones [69], scale up both model and dataset sizes by leveraging a mix of real and synthetic data, achieving strong generalization across diverse scenes. Although these methods have achieved significant progress on benchmark datasets, their accuracy still lags behind that of multi-view geometry-based approaches. In our system, we incorporate a monocular depth estimation module due to its robustness in low-texture regions and dynamic scenes to help improve overall performance and adaptability.
II-B Stereo matching
@@
Mainstream stereo matching methods recover disparity from matching costs. These methods can generally be divided into two categories: cost filtering-based methods and iterative optimization-based methods. Cost filtering-based methods [2, 22, 29, 70, 71, 9, 72] typically construct a 3D/4D cost volume using feature maps and subsequently employ 2D/3D CNN to filter the volume and derive the final disparity map. Constructing a cost volume with strong representational capacity and accurately regressing disparity from a noisy cost volume are key challenges. [1] proposes a stacked hourglass 3D CNN for better cost regularization. [2, 26, 73] propose the group-wise correlation volume, the attention concatenation volume, and the pyramid warping volume respectively to improve the expressiveness of the matching cost. Recently, a novel class of methods based on iterative optimization [4, 30, 6, 31, 74, 5] has achieved SOTA performance in both accuracy and efficiency. These methods use ConvGRUs to iteratively update the disparity by leveraging local cost values retrieved from the all-pairs correlation volume. Similarly, iterative optimization-based methods also primarily focus on improving the matching cost construction and the iterative optimization stages. CREStereo [30] proposes a cascaded recurrent network to update the disparity field in a coarse-to-fine manner. IGEV [6] proposes a Geometry Encoding Volume that encodes geometry and context information for a more robust matching cost. Both types of methods essentially derive depth from matching costs, which are inherently limited by ill-posed regions.
II-C Multi-View stereo
Numerous methods have been developed to estimate depth from multi-view images under the assumption of known camera intrinsics and extrinsics. Early work such as [75] pioneers the integration of deep feature learning into the multi-view stereo pipeline, although it still relies on traditional cost aggregation techniques. Later, [8] introduces a breakthrough approach by constructing a differentiable cost volume, followed by 3D convolutional neural networks (CNNs) for regularization, significantly improving depth estimation performance. Building upon this framework, [76] proposes a hybrid cost regularization network that decouples 3D local matching and 2D global context modeling to reduce computational overhead. Subsequent methods [77, 78, 79] further enhance performance and robustness through iterative residual refinement, while others [80, 81, 9] adopt coarse-to-fine strategies to balance accuracy and efficiency. More recently, transformer-based architectures [37, 82, 15] have been explored to enhance global context modeling for more robust feature representations. For example, MVSFormer [12] with hierarchical ViTs delivers substantial improvements over conventional pyramid convolution-based approaches. MVS2d[11] proposes an attention mechanism that aggregates features along the epipolar line for each query pixel in the reference image, significantly improving computational efficiency. Furthermore, MVSFormer++[15] introduces customized attention mechanisms tailored to different MVS modules, enabling better module adaptation. Despite these advances, the effectiveness of current MVS methods often hinges on scenes with large parallax and rich textures, and they struggle with dynamic content. Moreover, these approaches typically require precise relative pose estimation between views and are sensitive to errors in camera pose, limiting their robustness in practical applications.
II-D Multi-view depth estimation with Structural Priors
Multi-view depth estimation in ill-posed regions remains a significant challenge. To address this, prior work—both in stereo matching and multi-view stereo—has explored the integration of structural cues. For instance, EdgeStereo [35] incorporates edge detection features into the disparity estimation pipeline to enhance performance around object boundaries, while SegStereo [34] employs semantic segmentation cues to improve matching in textureless areas. However, such cues primarily offer object-level priors and lack the granularity required for accurate pixel-level depth estimation. As a result, these methods often struggle in scenes containing large curved or slanted surfaces. To provide finer structural guidance, several approaches have introduced monocular depth estimation as a complementary prior. CLStereo [83] incorporates a monocular branch to transfer geometric context into the stereo network, acting as a regularizer during matching. [84] significantly improves performance by fusing fine-grained local structures extracted from single-view methods with robust multi-view cues, particularly in regions with large parallax and strong gradients. [85] first predicts an initial depth map using a monocular depth estimation network, and then constructs a lightweight cost volume centered around the initial depth, thereby achieving a balance between accuracy and efficiency. Similarly, Los [86] utilizes monocular depth to generate local slanted planes that guide disparity refinement. FoundationStereo [87] employs a side-tuning feature backbone to adapt monocular priors from DepthAnythingV2 [38] to stereo inputs, thereby enhancing its generalization capability. AFNet [13] further enhances depth estimation performance and robustness through confidence-based adaptive fusion of predictions from the multi-view and single-view branches. In summary, all the aforementioned methods use monocular depth merely as prior. However, monocular depth is inherently affected by scale and shift ambiguities, as illustrated in Fig. 3, which limits its effective utilization and can introduce substantial noise when employed directly as a prior. To address this, our MonSter++ framework leverages monocular depth in a more robust manner by adaptively selecting reliable stereo disparities to calibrate the scale and shift of the monocular predictions. This design effectively integrates dense geometric priors while suppressing erroneous signals, leading to significantly improved performance in ill-posed regions.